# The HST-Hyperion Survey: Environmental Imprints on the Stellar-Mass Function at $z\sim 2.5$

### v1

@@ 3.3 Creating a Field Sample
To gauge the effect of overdense environments, we need a control set of galaxies that have grown in lower–density regions. We therefore draw two “field” slices that bracket Hyperion, $2.15\leq z\leq 2.25$ and $2.80\leq z\leq 2.90$. These ranges keep look-back-time differences small while minimizing the chance that Hyperion members scatter into the field sample (or vice-versa).
Defining the field solely as voxels with $n_{\sigma}<2.5$ would require excluding every voxel with $n_{\sigma}\geq 2.5$, erasing genuine large-scale fluctuations caused by cosmic variance and yielding an unrealistically uniform field. Instead, we remove only those galaxies that belong to a VMC-identified protostructure at redshifts $2\leq z\leq 3$, adopting the definition of a protostructure defined in the previous section.
@@ 4 Building the Stellar Mass Functions
Despite our extensive spectroscopic data within Hyperion, the total sample is still dominated by galaxies with relatively poorly constrained photo-$z$s (compared in Figures 1 and 2). In theory, photometric galaxies which fall outside of the range of Hyperion could really be located in one of its peaks, changing the SMF of Hyperion, and vice-versa. To extract as much information as we can from our sample we must be able to correctly quantify our uncertainty regarding which peak (if any) a galaxy is located in by correctly accounting for each galaxy’s redshift uncertainty. However, as the stellar masses are derived via SED-fitting, the redshift of a galaxy is intricately tied to the estimated stellar mass of the galaxy. Thus, a redshift uncertainty results not only in an uncertainty in the 3D location (and thus overdensity via the VMC maps) of a galaxy, but also in the galaxy’s stellar mass.
Rather than making assumptions about how the stellar mass changes as a function of redshift, we instead choose to sample the redshift distributions of our data via a Monte Carlo (MC) process, and re-fit the galaxy SED at each sampled redshift. In this section, we describe how we generate the MC realizations, and how each galaxy in a given realization is fit with a new stellar mass estimate generated in the process, before finally creating the SMFs based on the resulting galaxy distributions. It should be noted that this process is performed separately of the MC process used to generate the overdensity map (Section 3.1). In what follows, we sample only the redshifts of the galaxies and assign an overdensity based on the already-generated VMC map.
@@ 4.1 Probing Redshift Distributions
In order to account for redshift uncertainties, we create a suite of 100 Monte Carlo realizations in which we attempt to well-sample the redshift PDFs of our data. Broadly speaking, a given object in our sample will have a photo-$z$, spec-$z$, grism-$z$, or some combination of the three. The goal then is, for galaxies with more than just a photo-$z$, to use the more secure redshift measurement (either spec-$z$ or grism-$z$) a number of times which is proportional to the reliability measurement, given by the quality flag (see Section 2).
@@
For a given galaxy in a given MC iteration, the logic for choosing a redshift depends on the combination of redshift measurements available for that galaxy. For a galaxy with all three types of redshift measurements, we could potentially draw a new redshift from three different PDFs. Given each galaxy in our sample has COSMOS2020 photometry, they also have associated redshift PDFs from COSMOS2020 based on fits in LePhare. If the redshift in the given MC iteration is based on the photo-$z$, the redshift is drawn randomly from the COSMOS2020 PDF, $p(z)$. If based on the grism-$z$, redshifts are randomly drawn from a Gaussian distribution, $g(z)$, centered on the measured redshift from the grism spectroscopy, $z_{g}$, and with a width equal to $\sigma_{g}=46\cdot(1+z_{g})/14100$ (motivated in Section 2.3). Finally, if based on the spec-$z$, redshifts are simply taken to be the measured spectroscopic redshift, $z_{s}$. Thus, we outline how the random redshift, $z_{\rm MC}$, is drawn based on whether the galaxy has a photo-$z$ ($z_{p}$), spec-$z$ ($z_{s}$), grism-$z$ ($z_{g}$), or some combination of the three.
@@
Only $\mathbf{z_{p}}$: The vast majority of galaxies ($\sim 90.3\%$) only have a photo-$z$, which is reported in COSMOS2020. For these galaxies, we simply draw $z_{\rm MC}\sim p(z)$ for each MC iteration from the redshift PDF provided in COSMOS2020 from a LePhare fit.
@@
$\mathbf{z_{p}}$ and one of $\mathbf{z_{s}}$ or $\mathbf{z_{g}}$: In this case, the galaxy has COSMOS2020 photometry and either a $z_{g}$ or $z_{s}$. The spec-$z$ or grism-$z$ also has an associated reliability flag corresponding to some confidence in the redshift, $qf$ (described in Section 2). For each MC iteration, a random number, $\mathcal{R}$, is drawn from a uniform distribution on the half-open interval $[0,1)$. If $\mathcal{R}\geq qf$, then we take $z_{\rm MC}\sim p(z)$. Otherwise, we take $z_{\rm MC}=z_{s}$ or $z_{\rm MC}\sim g(z)$, depending on whether the additional redshift is a spec-$z$ or grism-$z$.
@@
$\mathbf{z_{p}},\,\,\mathbf{z_{s}}$ and $\mathbf{z_{g}}$: Some galaxies in our data have all three types of redshift measurements. In this case, the first step is to compare the confidence levels corresponding to the quality flags of the additional redshift measurements, which we call $qf_{\rm s}$ and $qf_{\rm g}$ for the spec-$z$ and grism-$z$, respectively. For clarity, we define the quantity $\mathcal{Q}\equiv qf_{\rm max}+qf_{\rm min}\cdot(1-qf_{\rm max})$, corresponding to a $qf_{\rm min}$-fraction of the way between $qf_{\rm max}$ and 1. That is, if $qf_{\rm max}=0.95$ and $qf_{\rm min}=0.7$, (corresponding to 95% and 70% confidences in the additional redshift measurements) then $\mathcal{Q}=0.95+0.7\cdot(0.05)=0.985$, or 70% of the way between 0.95 and 1. This way, if the less reliable of the additional redshifts is selected in a given MC iteration, the remaining interval $qf_{\rm max},1)$ is weighted according to the confidence associated with that less reliable measurement.
@@
With this defined, a random number, $\mathcal{R}$, is drawn from a uniform distribution on the half-open interval $[0,1)$, and a redshift is assigned depending on which of the additional redshift measurements has a higher confidence. In the case $qf_{\rm max}=qf_{\rm s}>qf_{\rm g}=qf_{\rm min}$ (the spec-$z$ is more reliable than the grism-$z$), a redshift is assigned according to:
$\begin{cases}\mathcal{R}<qf_{\rm max}&\Rightarrow z_{\rm MC}=z_{s}\\ qf_{\rm max}\leq\mathcal{R}<\mathcal{Q}&\Rightarrow z_{\rm MC}\sim g(z)\\ \mathcal{Q}\leq\mathcal{R}<1&\Rightarrow z_{\rm MC}\sim p(z)\end{cases}$
@@
Conversely, if $qf_{\rm max}=qf_{\rm g}>qf_{\rm s}=qf_{\rm min}$ (the grism-$z$ is more reliable than the spec-$z$), a redshift is assigned according to:
$\begin{cases}\mathcal{R}<qf_{\rm max}&\Rightarrow z_{\rm MC}\sim g(z)\\ qf_{\rm max}\leq\mathcal{R}<\mathcal{Q}&\Rightarrow z_{\rm MC}=z_{s}\\ \mathcal{Q}\leq\mathcal{R}<1&\Rightarrow z_{\rm MC}\sim p(z)\end{cases}$
@@
We apply this process to all $220,356$ galaxies in our sample, shown in Figure 1. In each MC iteration, we choose a redshift for all $220,356$ galaxies, of which $\sim 9.7\%$ additionally have a spectroscopic or grism redshift. For future steps in a given MC iteration we only consider galaxies which fall in the redshift range $2<z<3$.
@@ 4.2 Fitting Stellar Mass
For each Monte Carlo realization, we refit the stellar mass of every galaxy based on its MC-assigned redshift. In order to refit the masses, we use LePhare with similar parameters to those used in the COSMOS2015 and COSMOS2020 catalogs (Ilbert et al. 2015; Laigle et al. 2016; Weaver et al. 2022). Since we are working with high redshift data and are interested in a statistical study of a broad population of galaxies, we use only galaxy SED templates, excluding stellar and QSO templates. In this section, we briefly discuss the parameters used for the SED fits.
The SED fitting is performed using the flux densities reported in COSMOS2020. As described in Section 2.1, for a given a band XXXX, we use the aperture-to-total corrected flux density reported under XXXX_FLUX_APER3, and its error from the COSMOS2020 Classic catalog. When available, we use fluxes from 35 bands including all ground-based observations, Spitzer/IRAC channels 1 and 2, and NUV and FUV data from GALEX. We apply a systematic magnitude offset to the reported photometry (APPLY_SYSSHIFT in LePhare) given in Table 3 of Weaver et al. (2022).
@@
To ensure consistency of our fits with those reported in COSMOS2020, we refit a subset of relevant galaxies in the COSMOS2020 catalog and compare the resulting stellar masses. Specifically, we refit the $33,097$ galaxies which fall in the redshift range $2\leq z\leq 3$ in Figure 1, fixing the redshift to the value given by lp_zPDF in COSMOS2020. In Figure 4, we compare the stellar masses reported in COSMOS2020 via lp_mass_med with those from our refit. The fits are broadly consistent with 1:1, indicating the parameters used for SED fits in this study replicate those used in COSMOS2020 (Weaver et al. 2022) for the range of stellar masses we consider in this study.
Figure 4: Refitting COSMOS2020 Stellar Masses – We show the stellar masses of galaxies in the COMSMOS catalog (lp_MASS_MED) compared to our LePhare outputs as a consistency-check for out SED fitting parameters. The galaxies tested are those in Figure 1 in the redshift range $2\leq z\leq 3$ (as reported from lp_zPDF in COSMOS2020). The fits are consistent with the 1:1, indicating our SED fitting parameters are broadly consistent with those used in COSMOS2020.
@@ 4.3 Creating and Fitting the Stellar Mass Functions
We construct a stellar mass function for both Hyperion and the field in each of the 100 MC realizations. We restrict our analysis only to those galaxies which have $\log_{10}(M_{*}/M_{\odot})\geq 9.5$ to ensure a more complete sample. For each MC realization, we assign galaxies to either Hyperion or the field sample as defined in Sections 3.2 and 3.3, respectively. We then count the number of galaxies in each stellar mass bin, normalizing by the bin width and the volume of the corresponding population. We find no significant difference between the high- and low-redshift field samples, and therefore combine them into a single field SMF.
The volumes of Hyperion and its constituent overdensity peaks are computed as the sum of the volumes of the voxels representing them (see Section 3.1.1). For the field, we take the volume to be the rectangular prism that spans the full ranges in R.A., decl., and redshift for all galaxies, excluding the volumes of any protostructures within it (see Section 3.2).
@@
We additionally fit two forms of the Schechter function (Schechter 1976) to the Hyperion SMF. The single-Schechter form is given by:
$\begin{split}n_{\rm gal}&=\Phi\,\,d\log_{10}M\\ &=\ln(10)\Phi^{*}\,e^{-10^{\log_{10}M-\log_{10}M^{*}}}\\ &\quad\quad\times(10^{\log_{10}M-\log_{10}M^{*}})^{\alpha+1}\,\,d\log_{10}M.\end{split}$ (5)
@@
The best-fit Schechter-functions are shown in Figure 5, along with a single-Schechter fit to the COSMOS2020 field sample in the range $2.5<z\leq 3$, based on only photometry (Table C.1 in Weaver et al. 2023). Best-fit parameters and respective reduced $\chi^{2}$-values for the single- and double-Schechter fits are given in Table 3. We additionally report the Bayesian Information Criterion (BIC; Schwarz 1978), calculated via:
$\text{BIC}=k\ln(n)+\chi^{2}$ (7)
@@
In order to compare the SMF of Hyperion to that of the field, we perform a two-sample Kolmogorov-Smirnov (KS) test – a nonparametric method of testing whether two samples are drawn from the same underlying distribution. In this case, we treat each of the 100 MC realizations separately and perform a two-sample KS test comparing the stellar masses of the galaxies in the field to those in Hyperion. The resulting distribution of $p$-values for the 100 MC iterations is shown in the left panel of Figure 6, in which the null hypothesis is that the underlying mass distributions are the same for the field and Hyperion. In total, we find 33/100 MCs with a $p$-value <0.05<0.05.
@@
Figure 5: Hyperion Stellar Mass Function – We show the total SMF of Hyperion as well as the coeval field. Additionally, we show the single-Schechter fit of the COSMOS2020 fit for $2.5<z<3.0$ (magenta line; Weaver et al. 2023). Finally, we show the best-fit single- (green) and double-Schechter (blue) functions to the Hyperion SMF (see parameters in Table 3).
@@
Figure 6: Two Sample KS Tests – The results of two sample KS tests ran for each of the 100 MC realizations in which we compare the mass distribution of the field to that of (Left) Hyperion as a whole and (Right) Hyperion’s outskirts (pink) and overdensity peaks (orange). We find $33\%$ of MCs in which Hyperion as a whole is statistically different that the field ($p<0.05$), compared to $2\%$ and $97\%$ for the Hyperion outskirts and peaks, respectively. Thus, the underlying mass distribution for Hyperion’s peaks are significantly different from the field for most of the MC realizations, compared to Hyperion’s outskirts which show no significant difference.
@@
In addition to the SMF for Hyperion as a whole, we also build SMFs for different discrete overdensity thresholds within Hyperion. Specifically, we use the MC iterations to build SMFs for galaxies within the outskirts of Hyperion ($2.5\leq n_{\sigma}<4.0$) and the peaks of Hyperion ($4.0\leq n_{\sigma}$), where $n_{\sigma}$ is given by Equation 3. The resulting SMFs are shown in the left panel of Figure 7. For the field sample and each overdensity bin within Hyperion, we fit a single-Schechter function (Equation 5), allowing $\alpha$, $M^{*}$, and $\Phi^{*}$ to freely vary, as well as with a fixed $\alpha=-1.3$ – similar to methods adopted by a variety of studies (e.g., Marchesini et al. 2009; Muzzin et al. 2013; Tomczak et al. 2014). The fits are shown in the right panel of Figure 7, and the best-fit parameters as well as their respective reduced-$\chi^{2}$ and BIC values are given in Table 3. As before, the covariances of the parameters in the Schechter fits are discussed in Appendix B.
@@
We again perform a two-sample KS test to compare both the stellar mass distributions of Hyperion’s outskirts and peaks to that of the field. We treat each MC realization separately, and plot the resulting distribution of $p$-values in the right plot of Figure 6. When comparing mass distributions of Hyperion’s outskirts and the field, we find 2/100 MC realizations with a $p$-value <0.05<0.05. On the other hand, when comparing the peaks and the field, we find 97/100 MC realizations have a $p$-value <0.05<0.05.
@@
Figure 7: SMFs of Hyperion Peaks and Outskirts – (Left) The SMFs of the outskirts ($2.5\leq n_{\sigma}<4.0$) and overdensity peaks ($4.0\leq n_{\sigma}$) of Hyperion. The coeval field and COSMOS2020 fit are identical to Figure 5. (Right) The best-fit Schechter functions for the outskirts and overdensity peaks of Hyperion, and the coeval field. The solid lines correspond to single-Schechter fits allowing all three parameters to vary freely, and dot-dashed lines correspond to fits fixing the power-law slope to $\alpha=-1.3$. Best-fit parameters for each fit are given in Table 3.
@@
Empirically, trends appear approximately linear in log-log space. To quantify any trends, we fit the log-ratios of both the outskirts and peaks with a power-law of the form $\phi/\phi_{\rm field}=A\cdot(M_{*}/M_{\odot})^{\beta}$, where $A$ is a constant. This appears as a line with the form $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*})+\log_{10}(A)$ when plotted in log-log space. The best-fit slopes, $\beta$, and reduced $\chi^{2}$ values are reported in Table 1. A two-sided Student’s $t$-test performed during the linear regression analysis shows no significant correlation between stellar mass and the log-ratio in Hyperion’s outskirts relative to the field ($p=0.38$), but a strong correlation in the peaks ($p=0.02$). Despite the quality of the linear fits, one may expect the ratio of the SMFs to have a more complex shape given the relative complexity of the single-Schechter fits used to describe them. However, we show in Appendix C that a nearly-linear fit (in log-log space) is expected in the mass range considered in this work.
Table 1: Parameters for the SMF ratios – The power-law fits for the SMF-ratios ($\phi/\phi_{\rm field}$) shown in Figure 8. The columns are (1) the sample being fit, (2) the power-law slope of the fit, $\beta$, assuming $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*})+\log_{10}(A)$, and (3) the reduced-$\chi^{2}$ value of each fit.

### v2

@@ 3.3 Creating a field sample
To gauge the effect of overdense environments, we need a control set of galaxies that have grown in lower-density regions. We therefore drew from two field slices that bracket Hyperion, $2.15\leq z\leq 2.25$ and $2.80\leq z\leq 2.90$. These ranges kept look-back-time differences small while minimizing the chance that Hyperion members scattered into the field sample (or vice-versa).
Defining the field solely as voxels with $n_{\sigma}<2.5$ would require excluding every voxel with $n_{\sigma}\geq 2.5$, erasing genuine large-scale fluctuations caused by cosmic variance and yielding an unrealistically uniform field. Instead, we removed only those galaxies that belong to a VMC-identified protostructure at redshifts $2\leq z\leq 3$, adopting the definition of a protostructure defined in the previous section.
@@ 4 Building the stellar mass functions
Despite our extensive spectroscopic data within Hyperion, the total sample is still dominated by galaxies with relatively poorly constrained photo-$z$s (compared in Figures 1 and 2). In theory, photometric galaxies which fall outside of the range of Hyperion could in fact be located in one of its peaks, changing the SMF of Hyperion, and vice-versa. To extract as much information as we can from our sample we must correctly quantify our uncertainty regarding which peak (if any) a galaxy is located in by correctly accounting for each galaxy’s redshift uncertainty. However, as the stellar masses are derived via SED-fitting, the redshift of a galaxy is intricately tied to the estimated stellar mass of the galaxy. Thus, a redshift uncertainty results not only in an uncertainty in the 3D location (and thus overdensity via the VMC maps) of a galaxy, but also in the galaxy’s stellar mass.
Rather than making assumptions about how the stellar mass changes as a function of redshift, we instead chose to sample the redshift distributions of our data via a Monte Carlo (MC) process, and refit the galaxy SED at each sampled redshift. In this section, we describe how we generated the MC realizations, and how each galaxy in a given realization was fit with a new stellar mass estimate generated in the process, before finally creating the SMFs based on the resulting galaxy distributions. It should be noted that this process was performed separately from the MC process used to generate the overdensity map (Section 3.1). In what follows, we sampled only the redshifts of the galaxies and assigned an overdensity based on the already-generated VMC map.
@@ 4.1 Probing redshift distributions
In order to account for redshift uncertainties, we created a suite of 100 Monte Carlo realizations in which we attempted to accurately sample the redshift PDFs of our data. Broadly speaking, a given object in our sample has a photo-$z$, spec-$z$, grism-$z$, or some combination of the three. Therefore, the goal was, for galaxies with more than just a photo-$z$, to use the more secure redshift measurement (either spec-$z$ or grism-$z$) a number of times which is proportional to the reliability measurement, given by the quality flag (see Section 2).
@@
For a given galaxy in a given MC iteration, the logic for choosing a redshift depended on the combination of redshift measurements available for that galaxy. For a galaxy with all three types of redshift measurements, we potentially drew a new redshift from three different PDFs. Given each galaxy in our sample has COSMOS2020 photometry, they also have associated redshift PDFs from COSMOS2020 based on fits in LePhare. If the redshift in the given MC iteration was based on the photo-$z$, the redshift was drawn randomly from the COSMOS2020 PDF, $p(z)$. If based on the grism-$z$, redshifts were randomly drawn from a Gaussian distribution, $g(z)$, centered on the measured redshift from the grism spectroscopy, $z_{g}$, and with a width equal to $\sigma_{g}=46\cdot(1+z_{g})/14100$ (motivated in Section 2.3). Finally, if based on the spec-$z$, redshifts were simply taken to be the measured spectroscopic redshift, $z_{s}$. Thus, we outline how the random redshift, $z_{\rm MC}$, is drawn based on whether the galaxy has a photo-$z$ ($z_{p}$), spec-$z$ ($z_{s}$), grism-$z$ ($z_{g}$), or some combination of the three.
Only $z_{p}$
@@
: The vast majority of galaxies ($\sim 90.3\%$) only have a photo-$z$, which is reported in COSMOS2020. For these galaxies, we simply drew $z_{\rm MC}\sim p(z)$ for each MC iteration from the redshift PDF provided in COSMOS2020 from a LePhare fit.
$z_{p}$ and one of $z_{s}$ or $z_{g}$
@@
: In this case, the galaxy has COSMOS2020 photometry and either a $z_{g}$ or $z_{s}$. The spec-$z$ or grism-$z$ also has an associated reliability flag corresponding to some confidence in the redshift, $qf$ (described in Section 2). For each MC iteration, a random number, $\mathcal{R}$, was drawn from a uniform distribution on the half-open interval $[0,1)$. If $\mathcal{R}\geq qf$, then we took $z_{\rm MC}\sim p(z)$. Otherwise, we took $z_{\rm MC}=z_{s}$ or $z_{\rm MC}\sim g(z)$, depending on whether the additional redshift is a spec-$z$ or grism-$z$.
$z_{p},\,\,z_{s}$ and $z_{g}$
@@
: Some galaxies in our data have all three types of redshift measurements. In this case, the first step was to compare the confidence levels corresponding to the quality flags of the additional redshift measurements, which we call $qf_{\rm s}$ and $qf_{\rm g}$ for the spec-$z$ and grism-$z$, respectively. For clarity, we define the quantity $\mathcal{Q}\equiv qf_{\rm max}+qf_{\rm min}\cdot(1-qf_{\rm max})$, corresponding to a $qf_{\rm min}$-fraction of the way between $qf_{\rm max}$ and 1. That is, if $qf_{\rm max}=0.95$ and $qf_{\rm min}=0.7$, (corresponding to 95% and 70% confidences in the additional redshift measurements) then $\mathcal{Q}=0.95+0.7\cdot(0.05)=0.985$, or 70% of the way between 0.95 and 1. This way, if the less reliable of the additional redshifts is selected in a given MC iteration, the remaining interval $qf_{\rm max},1)$ is weighted according to the confidence associated with that less reliable measurement.
@@
With this defined, a random number, $\mathcal{R}$, was drawn from a uniform distribution on the half-open interval $[0,1)$, and a redshift was assigned depending on which of the additional redshift measurements had a higher confidence. In the case $qf_{\rm max}=qf_{\rm s}>qf_{\rm g}=qf_{\rm min}$ (the spec-$z$ is more reliable than the grism-$z$), a redshift was assigned according to
$\begin{cases}\mathcal{R}<qf_{\rm max}&\Rightarrow z_{\rm MC}=z_{s}\\ qf_{\rm max}\leq\mathcal{R}<\mathcal{Q}&\Rightarrow z_{\rm MC}\sim g(z)\\ \mathcal{Q}\leq\mathcal{R}<1&\Rightarrow z_{\rm MC}\sim p(z)\end{cases}.$
@@
Conversely, if $qf_{\rm max}=qf_{\rm g}>qf_{\rm s}=qf_{\rm min}$ (the grism-$z$ was more reliable than the spec-$z$), a redshift was assigned according to
$\begin{cases}\mathcal{R}<qf_{\rm max}&\Rightarrow z_{\rm MC}\sim g(z)\\ qf_{\rm max}\leq\mathcal{R}<\mathcal{Q}&\Rightarrow z_{\rm MC}=z_{s}\\ \mathcal{Q}\leq\mathcal{R}<1&\Rightarrow z_{\rm MC}\sim p(z)\end{cases}.$
@@
We applied this process to all $220,356$ galaxies in our sample, shown in Figure 1. In each MC iteration, we chose a redshift for all $220,356$ galaxies, of which $\sim 9.7\%$ additionally had a spectroscopic or grism redshift. For future steps in a given MC iteration we only considered galaxies which fell in the redshift range $2<z<3$.
@@ 4.2 Fitting stellar mass
For each Monte Carlo realization, we refitted the stellar mass of every galaxy based on its MC-assigned redshift. In order to refit the masses, we used LePhare with similar parameters to those used in the COSMOS2015 and COSMOS2020 catalogs (Ilbert et al. 2015; Laigle et al. 2016; Weaver et al. 2022). Since we are working with high-redshift data and are interested in a statistical study of a broad population of galaxies, we used only galaxy SED templates, excluding stellar and QSO templates. In this section, we briefly discuss the parameters used for the SED fits.
The SED fitting was performed using the flux densities reported in COSMOS2020. As described in Section 2.1, for a given a band XXXX, we used the aperture-to-total corrected flux density reported under XXXX_FLUX_APER3, and its error from the COSMOS2020 Classic catalog. When available, we used fluxes from 35 bands including all ground-based observations, Spitzer/IRAC channels 1 and 2, and NUV and FUV data from GALEX. We applied a systematic magnitude offset to the reported photometry (APPLY_SYSSHIFT in LePhare) given in Table 3 of Weaver et al. (2022).
@@
To ensure consistency of our fits with those reported in COSMOS2020, we refitted a subset of relevant galaxies in the COSMOS2020 catalog and compared the resulting stellar masses. Specifically, we refitted the $33,097$ galaxies which fall in the redshift range $2\leq z\leq 3$ in Figure 1, fixing the redshift to the value given by lp_zPDF in COSMOS2020. In Figure 5, we compare the stellar masses reported in COSMOS2020 via lp_mass_med with those from our refit. The fits are broadly consistent with 1:1, indicating the parameters used for SED fits in this study replicate those used in COSMOS2020 (Weaver et al. 2022) for the range of stellar masses we consider in this study.
Figure 5: We show the stellar masses of galaxies in the COMSMOS catalog (lp_MASS_MED) compared to our LePhare outputs as a consistency check for our SED fitting parameters. The galaxies tested are those in Figure 1 in the redshift range $2\leq z\leq 3$ (as reported from lp_zPDF in COSMOS2020). The fits are consistent with the 1:1, indicating our SED fitting parameters are broadly consistent with those used in COSMOS2020.
@@ 4.3 Creating and fitting the stellar mass functions
We constructed a stellar mass function for both Hyperion and the field in each of the 100 MC realizations. We restricted our analysis only to those galaxies which have $\log_{10}(M_{*}/M_{\odot})\geq 9.5$ to ensure a more complete sample. For each MC realization, we assigned galaxies to either Hyperion or the field sample as defined in Sections 3.2 and 3.3, respectively. We then counted the number of galaxies in each stellar mass bin, normalizing by the bin width and the volume of the corresponding population. We find no significant difference between the high- and low-redshift field samples, and therefore combined them into a single field SMF.
The volumes of Hyperion and its constituent overdensity peaks were computed as the sum of the volumes of the voxels representing them (see Section 3.1.1). For the field, we took the volume to be the rectangular prism that spans the full ranges in R.A., Dec., and redshift for all galaxies, excluding the volumes of any protostructures within it (see Section 3.2).
@@
We additionally fitted two forms of the Schechter function (Schechter 1976) to the Hyperion SMF. The single-Schechter form is given by
$\begin{split}n_{\rm gal}&=\Phi\,\,d\log_{10}M\\ &=\ln(10)\Phi^{*}\,e^{-10^{\log_{10}M-\log_{10}M^{*}}}\\ &\quad\quad\times(10^{\log_{10}M-\log_{10}M^{*}})^{\alpha+1}\,\,d\log_{10}M,\end{split}$ (5)
@@
The best-fit Schechter-functions are shown in Figure 6, along with a single-Schechter fit to the COSMOS2020 field sample in the range $2.5<z\leq 3$, based on only photometry (Table C.1 in Weaver et al. 2023). Best-fit parameters and respective reduced $\chi^{2}$-values for the single- and double-Schechter fits are given in Table 3. We additionally report the Bayesian information criterion (BIC; Schwarz 1978), calculated via
$\text{BIC}=k\ln(n)+\chi^{2},$ (7)
@@
In order to compare the SMF of Hyperion to that of the field, we performed a two-sample Kolmogorov-Smirnov (KS) test – a nonparametric method of testing whether two samples are drawn from the same underlying distribution. In this case, we treated each of the 100 MC realizations separately and performed a two-sample KS test comparing the stellar masses of the galaxies in the field to those in Hyperion. The resulting distribution of $p$-values for the 100 MC iterations is shown in the left panel of Figure 7, in which the null hypothesis is that the underlying mass distributions are the same for the field and Hyperion. In total, we find 33/100 MCs with a $p$-value <0.05<0.05.
@@
Figure 6: We show the total SMF of Hyperion as well as the coeval field. Additionally, we show the single-Schechter fit of the COSMOS2020 fit for $2.5<z<3.0$ (magenta line; Weaver et al. 2023). Finally, we show the best-fit single-Schechter (green) and double-Schechter (blue) functions to the Hyperion SMF (see parameters in Table 3).
@@
Figure 7: The results of two sample KS tests that were run for each of the 100 MC realizations in which we compare the mass distribution of the field to that of (left) Hyperion as a whole and (right) Hyperion’s outskirts (pink) and overdensity peaks (orange). We find $33\%$ of MCs in which Hyperion as a whole is statistically different from that of the field ($p<0.05$), compared to $2\%$ and $97\%$ for the Hyperion outskirts and peaks, respectively. Thus, the underlying mass distribution for Hyperion’s peaks are significantly different from the field for most of the MC realizations, compared to Hyperion’s outskirts, which show no significant difference.
@@
In addition to the SMF for Hyperion as a whole, we also built SMFs for different discrete overdensity thresholds within Hyperion. Specifically, we used the MC iterations to build SMFs for galaxies within the outskirts of Hyperion ($2.5\leq n_{\sigma}<4.0$) and the peaks of Hyperion ($4.0\leq n_{\sigma}$), where $n_{\sigma}$ is given by Equation 3. The resulting SMFs are shown in the left panel of Figure 8. For the field sample and each overdensity bin within Hyperion, we fitted a single-Schechter function (Equation 5), allowing $\alpha$, $M^{*}$, and $\Phi^{*}$ to freely vary, as well as with a fixed $\alpha=-1.3$ – similar to methods adopted by a variety of studies (e.g., Marchesini et al. 2009; Muzzin et al. 2013; Tomczak et al. 2014). The fits are shown in the right panel of Figure 8, and the best-fit parameters as well as their respective reduced-$\chi^{2}$ and BIC values are given in Table 3. As before, the covariances of the parameters in the Schechter fits are discussed in Appendix B.
@@
We again performed a two-sample KS test to compare both the stellar mass distributions of Hyperion’s outskirts and peaks to that of the field. We treated each MC realization separately, and plot the resulting distribution of $p$-values in the right plot of Figure 7. When comparing mass distributions of Hyperion’s outskirts and the field, we find 2/100 MC realizations with a $p$-value <0.05<0.05. On the other hand, when comparing the peaks and the field, we find 97/100 MC realizations have a $p$-value <0.05<0.05.
@@
Figure 8: Left: SMFs of the outskirts ($2.5\leq n_{\sigma}<4.0$) and overdensity peaks ($4.0\leq n_{\sigma}$) of Hyperion. The coeval field and COSMOS2020 fit are identical to Figure 6. Right: Best-fit Schechter functions for the outskirts and overdensity peaks of Hyperion, and the coeval field. The solid lines correspond to single-Schechter fits allowing all three parameters to vary freely; the dot-dashed lines correspond to fits fixing the power-law slope to $\alpha=-1.3$. The best-fit parameters for each fit are given in Table 3.
@@
Empirically, trends appear approximately linear in log-log space. To quantify any trends, we fitted the log-ratios of both the outskirts and peaks with a power-law of the form $\phi/\phi_{\rm field}=A\cdot(M_{*}/M_{\odot})^{\beta}$, where $A$ is a constant. This appears as a line with the form $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*})+\log_{10}(A)$ when plotted in log-log space. The best-fit slopes, $\beta$, and reduced $\chi^{2}$ values are reported in Table 1. A two-sided Student’s $t$-test performed during the linear regression analysis shows no significant correlation between stellar mass and the log-ratio in Hyperion’s outskirts relative to the field ($p=0.38$), but a strong correlation in the peaks ($p=0.02$). Despite the quality of the linear fits, one may expect the ratio of the SMFs to have a more complex shape given the relative complexity of the single-Schechter fits used to describe them. However, we show in Appendix C that a nearly linear fit (in log-log space) is expected in the mass range considered in this work.
To assess the robustness of the above results to uncertainties inherent to photometric redshifts, we repeated the SMF analysis for only members of Hyperion with either a ground-based or HST/grism spectroscopic sample. We find the difference between the SMFs of the overdensity peaks and the field is effectively unchanged, owing to the large fraction of galaxies in the peaks which have spectroscopic redshifts ($\sim 68/72$ in each MC iteration). The lack of difference between the shape of the SMF of Hyperion’s outskirts and that of the field is also effectively unchanged, though the normalization of the SMF of the outskirts is lowered due to the smaller fraction of spectroscopic redshifts ($\sim 90/134$ in each MC iteration). Therefore, the results do not appear to be driven by the inclusion of galaxies with only photometric redshifts, or the scattering of such galaxies into and out of Hyperion in the Monte Carlo process.
We also tested whether the SMF of overdensity peaks depends on large-scale substructure within Hyperion by subdividing the peak population according to redshift and right ascension. In particular, we find that the eastern ($\alpha\geq 150.2^{\circ}$) and western ($\alpha<150.52^{\circ}$) portions of the lower-$z$ region ($z<2.52$) of Hyperion have similar SMFs. Because the most massive peak, Theia, is entirely contained within the western, lower-$z$ region (Table 2), this agreement demonstrates that the observed SMF differences are not exclusive to the most massive and overdense region of Hyperion, but instead reflect a more general dependence on local overdensity.
Table 1: Parameters for the SMF ratios. [^2] The power-law fits for the SMF ratios ($\phi/\phi_{\rm field}$) shown in Figure 9. The columns are (1) the sample being fit; (2) the power-law slope of the fit, $\beta$, assuming $\log_{10}(\phi/\phi_{\rm field})=\beta\log_{10}(M_{*})+\log_{10}(A)$; and (3) the reduced-$\chi^{2}$ value of each fit.